# Token Communications: A Unified Framework for Cross-modal Context-aware Semantic Communications

### v1

@@ VI-B Proposed Cross-Modality TokCom Scheme
To enhance the TCE of TokCom under various channel conditions, we propose a TokCom w/ CMI scheme, which aims to reduce the overhead of retransmissions and predict lost token packets by exploiting the context and cross-modality information (CMI). In this scheme, each packet is transmitted only once, resulting in a retransmission count of $T=1$. This eliminates the need for packet re-transmissions, and hence, improves the TCE. For packets that experience errors, the corresponding token positions are marked with a special token, [MASK], in the decoded token sequence. The decoder then iteratively predicts these masked tokens by leveraging the contextual information surrounding them, using a pre-trained bi-directional transformer model called MaskGIT [7]. To further improve the prediction accuracy of the lost tokens, we incorporate CMI as a conditioning signal. Specifically, a 7-bit image class label, which is transmitted through a separate channel, is assumed to be perfectly received by the receiver. This label provides additional contextual information that helps the transformer make more accurate predictions for the lost tokens, improving quality of the received image.
In addition to the proposed “TokCom w/ CMI” scheme, we also consider two baseline comparison schemes to evaluate the performance improvements. Conventional scheme: In this scheme, the receiver requests a retransmission for any errored packet, ensuring that all tokens are received correctly. For simplicity in analysis, the signaling overhead to request retransmissions over the feedback channel is not considered; however, including this overhead would further decrease the overall efficiency. TokCom w/o CMI: This baseline version of TokCom operates without the use of CMI, providing a simpler alternative. It serves as a reference for assessing the impact of incorporating CMI on the system performance.
Figure 4: Cross-modality TokCom performance for generative image semantic communication. The metric TCE represents the token communication bandwidth efficiency. The perceptual quality is measured by CLIP (↑) and LPIPS (↓), and the pixel distortions are measured by PSNR (↑).
@@ VI-C Performance Analysis
As shown in Fig. 4, the proposed TokCom framework significantly enhances performance across different channel conditions. At an SNR of $10$ dB, where the packet error rate (PER) is moderate at $19\%$, the system achieves a substantial $23.8\%$ improvement in TCE, while maintaining a CLIP score comparable with the conventional baseline. This demonstrates that TokCom not only improves efficiency but also preserves the semantic quality of the transmitted image. When the channel conditions degrade to an SNR of $6$ dB, resulting in a higher PER of $41\%$, TokCom exhibits remarkable robustness. The system experiences only a slight $4.5\%$ degradation in CLIP score, maintaining a score above $0.7$, which highlights its ability to handle more challenging communication environments without significant performance loss.
Furthermore, a comparison between the two TokCom variants reveals that the “TokCom w/o CMI” version performs worse, with lower CLIP scores, when compared to “TokCom w/ CMI”. This empirically validates the importance of incorporating CMI in preserving semantic quality. As shown in the visual samples of Fig. 4, the class label “Pit Bulls” serves as CMI to guide the lost image token prediction, enabling the receiver to reconstruct a Pit Bull with high perceptual quality, even with a large PER. The performance gap between these variants underscores the vital role CMI plays in improving prediction accuracy and preserving image semantics.
Additionally, as shown in Fig. 4, the generative token prediction capability of TokCom ensures that both learned perceptual image patch similarity (LPIPS) and peak signal-to-noise ratio (PSNR) degrade gradually with increasing PER, similar to the trend observed in CLIP. This demonstrates that TokCom consistently preserves high perceptual quality even under challenging error conditions. Finally, while incorporating CMI leads to improvements in semantic and perceptual quality metrics, i.e. CLIP and LPIPS, its impact on distortion metrics like PSNR is minimal. This underscores the importance of semantic and perceptual quality metrics in evaluating generative semantic communication, as they more accurately reflect the quality of reconstructed content in this context.
@@ VII-B TokCom Computational Complexity and Collaborative Inference
According to the scaling laws, performance of LMs reliably improve with their size and computational complexity [8]. The existing GFM/MLLMs are typically computationally complex, thereby posing challenges in their deployment in the TokCom framework. To tackle this, a collaborative device-edge-cloud inference should be designed, where lightweight models can run on-device, while more complex inference is offloaded to edge or cloud servers equipped with high-performance larger models. The offloading strategy should optimize the trade off between computational complexity, latency, and the resulting semantic quality. The corresponding latency-performance tradeoffs should be derived for TokCom by considering factors such as model size, computational resources, channel conditions, and the heterogeneity in zero/few-shot performance [14] of pre-trained MLLMs deployed on device-edge-cloud. As an example, [15] optimized the offloading strategy with a matching approach to minimize the total communication/computation latency while maximizing the semantic quality.
@@ VII-C TokCom Privacy and Security
The use of tokens as communication units and the reliance on pre-trained GFM/LLMs introduces potential vulnerabilities for TokCom. Semantic payloads are vulnerable to inference attacks, necessitating robust encryption and privacy-preserving mechanisms. Adversaries could potentially intercept or manipulate token sequences, leading to unauthorized access to sensitive information or alteration of transmitted content. Additionally, the use of context information in TokCom systems may inadvertently reveal more about the communicating parties than intended. To address these challenges, robust encryption methods specifically designed for token-based communications need to be developed. It is also crucial to consider the security implications of using pre-trained GFM/MLLMs, as they may be susceptible to adversarial attacks or contain biases that could compromise the integrity of the TokCom system. As TokCom evolves, ensuring privacy and security at both the architectural and operational levels will remain paramount to its effective deployment in future wireless networks.

### v3

@@ VI-B Proposed Cross-Modality TokCom Scheme
To enhance the TCE of TokCom under various channel conditions, we propose a “TokCom w/ CMI” scheme, which aims to reduce the overhead of retransmissions and predict lost token packets by exploiting the context and cross-modality information (CMI). In this scheme, each packet is transmitted only once, resulting in a retransmission count of $T=1$. This eliminates the need for packet re-transmissions, and hence, improves the TCE. For packets that experience errors, the corresponding token positions are marked with a special token, [MASK], in the decoded token sequence. The decoder then iteratively predicts these masked tokens by leveraging the contextual information surrounding them, using a pre-trained bi-directional transformer model called MaskGIT [7] with the VQGAN-based tokenizer. To further improve the prediction accuracy of the lost tokens, we incorporate CMI as a conditioning signal. Specifically, a 7-bit image class label, which is transmitted through a separate channel, is assumed to be perfectly recovered by the receiver and used as CMI. This label provides additional contextual information that helps the transformer make more accurate predictions for the lost tokens, improving quality of the received image.
In addition to the proposed “TokCom w/ CMI” framework, we evaluate several baseline schemes to assess the performance improvements achieved by TokCom. As TokCom enables digital semantic communication, we have used digital image compression/SemCom benchmarks working at similar bpp values, and all schemes use identical settings for packet length, channel coding, and modulation for a fair comparison. “Cheng [11]” adopts the recently developed powerful learned image compression method [11], with the pre-trained model fine-tuned to achieve a bpp of $\approx 0.04$. “VQGAN [12]” employs the tokenizer but without transformer, to demonstrate the gains from transformer-based context processing in TokCom. We also simulate “Cheng [11] + R” and “VQGAN [12] + R”, in which the receiver requests retransmissions for any corrupted packets to guarantee accurate reception of all tokens. Finally, “TokCom w/o CMI” represents a simplified version of TokCom that omits the CMI, serving as a baseline to quantify the performance gains due to the use of CMI.
Figure 4: Cross-modality TokCom performance for generative image semantic communication. The TCE metric represents the token communication bandwidth efficiency. The semantic and perceptual quality are measured by CLIP (↑) and LPIPS (↓) metrics, while the pixel distortions are measured by PSNR (↑).
@@ VI-C Performance Analysis
In Fig. 4 we have compared the proposed TokCom framework with and without cross modality information (CMI) with the above benchmarks, and the results demonstrate significant gains for TokCom. Firstly, the conventional scheme, i.e., Cheng [11] benchmark, fails to achieve acceptable reconstruction quality without retransmissions in presence of higher packet errors, i.e., the PSNR for this benchmark drops below 10dB when PER>~{}>~{}>0.07 or SNR${}<8$ dB, which leads to unacceptable visual quality. On the other hand, the VQGAN [12] benchmark provides the performance closer to that of TokCom, but the performance gap in the quality metrics is considerable as the PER grows, specifically in terms of the CLIP metric that demonstrates the semantic quality. This highlights the gains from transformer-based context processing that enables mitigating errors without requiring retransmissions. On the other hand, when we apply retransmissions in the “Cheng [11] + R” and “VQ-GAN [12] + R” benchmarks, the TCE drops severely specifically for higher PER values. Notably, the “Cheng [11] + R” scheme generates significantly blurrier images under the same low bpp setting and delivers inferior semantic and perceptual quality compared to TokCom, as evidenced by both CLIP and LPIPS metrics. Another interesting observation is that VQ-GAN [12] performs better than Cheng [11] in terms of the CLIP and LPIPS metrics which better demonstrate the semantic/perceptual quality, although Cheng [11] performs better in terms of the PSNR which is more of a distortion metric and of less importance in SemCom.
Furthermore, a comparison between the two TokCom variants reveals that the “TokCom w/o CMI” version performs worse, with lower CLIP scores, when compared to “TokCom w/ CMI”. This validates the importance of incorporating CMI in preserving semantic quality. As shown in the visual samples of Fig. 4, the class label “Pit Bulls” serves as CMI to guide the lost image token prediction, enabling the receiver to reconstruct a Pit Bull with high perceptual quality, even with a large PER. The performance gap between these variants underscores the vital role that CMI plays in improving prediction accuracy and preserving image semantics. We specifically see that leveraging cross-modal context information to avoid retransmissions increases the bandwidth efficiency with negligible loss of semantic/perceptual quality. At a moderate PER of $20\%$, the system achieves a substantial $24.4\%$ improvement in TCE, while maintaining a CLIP score comparable with the conventional baseline. This demonstrates that TokCom not only improves bandwidth efficiency but also preserves the semantic quality of the transmitted image. When the channel conditions degrade to an SNR of $6$ dB, resulting in a higher PER of $40\%$, TokCom exhibits remarkable robustness. The system experiences only a slight $4.2\%$ degradation in CLIP score, maintaining a score above $0.7$, which highlights its ability to handle more challenging communication environments without significant semantic loss. Additionally, as shown in Fig. 4, the generative token prediction capability of TokCom ensures that both LPIPS and PSNR degrade gradually with increasing PER, similar to the trend observed in CLIP. This demonstrates that TokCom consistently preserves high perceptual quality even under challenging error conditions. Finally, while incorporating CMI leads to improvements in semantic and perceptual quality metrics, i.e., CLIP and LPIPS, its impact on distortion metrics like PSNR is minimal. This underscores the importance of semantic and perceptual quality metrics in evaluating generative semantic communication, as they more accurately reflect the quality of reconstructed content in this context. Finally, the received TER values before and after transformer-based context processing are [0.15, 0.1, 0.065, 0.038, 0.02, 0.005] and [0.03, 0.025, 0.019, 0.0126, 0.00764, 0.0021], respectively, corresponding to SNRs of [6, 6.5, 7, 7.5, 8, 9] dB. The result confirms effectiveness of transformer-based context processing in reducing the TER.
In terms of the communication/computation latency tradeoffs, we have used the VQGAN and MaskGIT for tokenization and context processing in simulations of Fig. 4, which requires a total of $\approx 0.8$ TFLOPs per image. Considering two recent edge AI chips, NVIDIA DGX Spark and Jetson AGX Orin with specifications 1000 FP4 TOPS and 275 INT8 TOPS (1 TFLOP (FP32) $\approx$ 12 TOPS (INT8) $\approx$ 24 TOPS (FP4)), the computation time per image is 19.2 and 34.9 msec, respectively. On the other hand, considering the bpp $\approx 0.04$ and the modulation/coding used, as well as the average number of retransmissions for each SNR/PER value, the communication time per image is [43.7, 36.9, 32.8, 29.8, 28.2, 26.7] msec for the SNRs [6, 6.5, 7, 7.5, 8, 9] dB, respectively. Considering the rapid growth in AI computation hardware, the above results provide strong evidence that additional computation time should be spent to run the large models in TokCom in order to avoid retransmissions and save on the communication time in more adverse channel/network conditions when the PER is high.
VI-D Other Modalities, Datasets, and Tokenizers
Fig. 5 presents examples of TokCom on Flicker image [^7] https://www.kaggle.com/datasets/hsankesara/flickr-image-dataset and ESC-50 audio [^8] https://github.com/karolpiczak/ESC-50 datasets incorporating audio and text CMI using another SOTA tokenizer and transformer context processing model, i.e., TiTok [^9] https://github.com/bytedance/1d-tokenizer/blob/main/README_MaskGen.md. The CLAP and CLIP models are used for audio and text CMI, respectively. The token codebook size is $Q=8192$, and each image is represented by 128 tokens, resulting in a bitrate of 0.025 bpp. At the receiver side, masked token prediction modeling is used to leverage the CMI for recovering corrupted tokens. These results show applicability of the proposed TokCom framework to other modalities, datasets, and tokenizer models.
Figure 5: Illustration of TokCom performance with audio and text-based CMI.
@@ VII-B TokCom Computational Complexity and Collaborative Inference
According to the scaling laws, performance of LMs reliably improve with their size and computational complexity [6]. The existing GFM/MLLMs are typically computationally complex, thereby posing challenges in their deployment in the TokCom framework. To tackle this, future research could explore design of collaborative device-edge-cloud inference schemes, where lightweight models can run on-device, while more complex inference is offloaded to edge or cloud servers equipped with high-performance larger models. The offloading strategy should optimize the trade off between computational complexity, latency, and the resulting semantic quality. The corresponding latency-performance tradeoffs should be derived for TokCom by considering factors such as model size, computational resources, channel conditions, and the heterogeneity in zero/few-shot performance of pre-trained GFM/MLLMs deployed on device-edge-cloud [13]. Another solution could be reasoning through collaboration between large and small models leveraging speculative decoding techniques [14], which could be further studied in future research.
@@ VII-C TokCom Privacy and Security
The use of tokens as communication units and the reliance on large pre-trained GFM/LLMs introduces potential vulnerabilities. Tokenized representations carry semantic payloads, and hence are more vulnerable to inference attacks. An adversary intercepting a TokCom flow can easier manipulate the transmitted content through tokens. The strong context information in token sequences may lead to leakage of sensitive information, e.g., about the communication intent. Future work could explore design of new encryption and privacy-preservation techniques in the token embedding space to mitigate these challenges. Moreover, using pre-trained GFM/MLLMs could be susceptible to adversarial attacks. Differential privacy schemes should be explored for pre-training to prevent leakage of any sensitive information from the training data. Pre-trained models may inherit or amplify biases present in their training data, which could adversely affect TokCom systems. The use of machine unlearning techniques could be explored to remove influence of any biased training data. As TokCom evolves, ensuring privacy and security at both the architectural and operational levels will remain paramount to its effective deployment in future wireless networks.
VII-D 6G Applications and Network Architecture
Finally, future research could explore emerging 6G applications such as the wireless metaverse and immersive communications, extended/mixed reality (XR/MR), the internet of senses, and holographic teleportation, where TokCom can enable efficient intent-aware communications leveraging cross-modal context information. To enable this, design of semantic-aware network architecture should be explored to allow distributed GFM/MLLMs deployment and support dynamic Knowledge Base (KB) management across the network layers. Scalable resource allocation schemes for TokCom on the semantic-aware radio access networks [15] should be developed to adapt not only to the channel conditions and device capabilities, but also to the multimodal context information.